# Finite Groups of Symplectic Automorphisms of Supersingular K3 surfaces in Odd Characteristics

### v2

@@ Abstract.
In 2009, Dolgachev–Keum show that finite groups of tame symplectic automorphisms of K3 surfaces in positive characteristics are subgroups of the Mathieu group of degree $23$ under mild conditions. In this paper, we utilize lattice-theoretic methods to investigate symplectic actions of finite groups $G$ on K3 surfaces in odd characteristics. In the tame cases (i.e., the order of $G$ is coprime with $p$) and all the superspecial cases, one can associate with the action a Leech pair which can be detected via Höhn–Mason’s list. Notice that the superspecial cases have been resolved by Ohashi–Schütt. If the K3 surface is supersingular with Artin invariant at least two, we develop a new machinery called $p$-root pairs to detect possible symplectic finite group actions (without the assumption of tameness). The concept of $p$-root pair is closely related to root systems and Weyl groups. In particular, we recover many results by Dolgachev–Keum with a different method and give an upper bound for the exponent of $p$ in $|G|$.
@@ 1. Introduction
K3 surfaces play an important role in many areas of mathematics and the study of their automorphism groups has a long and rich history. Nikulin [Nik79b] first systematically studied automorphisms of K3 surfaces via lattice theory. After that, in his pioneering work [Muk88], Mukai proved (using the so-called Mathieu representations) that a finite group acts faithfully and symplectically on a complex K3 surface if and only if it is a subgroup of the Mathieu group $M_{23}$ with the induced action on the set of $24$ points having at least $5$ orbits. The problem of classifying (symplectic) automorphism groups of K3 surfaces over fields of positive characteristics turns out to be more complicated. In [DK09], Dolgachev and Keum extended Mukai’s methods to classify finite groups $G$ that act symplectically and faithfully on K3 surfaces over an algebraically closed field of characteristic $p>0$ under the tameness condition that $\gcd(|G|,p)=1$. They also showed that if $p\geq 13$, then such $G$ has to be tame. Ohashi–Schütt [OS24] recently completely classified maximal finite groups of symplectic automorphisms of superspecial K3 surfaces in any positive characteristic.
In this paper, we investigate finite groups acting symplectically on K3 surfaces in positive characteristic $p\geq 3$ via the induced action of the group on the Néron–Severi lattice. Our method can be applied to group actions without the tameness assumption. Before stating our main results, we fix some terminologies.
Throughout this paper, let $p$ be an odd prime, and $k$ be any algebraically closed field with $\operatorname{char}(k)=p$. Let $X$ be a K3 surface in characteristic $p$ (which means that it is a K3 surface over $k$ in this paper). We denote the Néron–Severi lattice of $X$ by $\operatorname{NS}(X)$ with rank $\rho(X)$ (which is the Picard number of $X$). It is known that $\operatorname{NS}(X)$ is an even lattice with signature $(1,\rho(X)-1)$. The discriminant group of the lattice $\operatorname{NS}(X)$ is denoted by $A_{\operatorname{NS}(X)}$. We denote the automorphism group of $X$ by $\operatorname{Aut}(X)$. An automorphism $g$ of $X$ is symplectic if and only if $g$ acts identically on $H^{0}(X,\omega_{X})$. Here $\omega_{X}$ is the canonical line bundle on $X$. We denote by $\operatorname{Aut}^{s}(X)$ the group of symplectic automorphisms of $X$. Given a finite subgroup $G\leq\operatorname{Aut}^{s}(X)$, there is an induced action of $G$ on the Néron–Severi lattice $\operatorname{NS}(X)$.
@@
Suppose a group $G$ acts faithfully on a lattice $L$ such that $L^{G}=\{0\}$ and $G$ acts trivially on $A_{L}$. Then we denote by $\widetilde{G}$ the group of all automorphisms of $L$ which act trivially on $A_{L}$. We call $\widetilde{G}$ and $(\widetilde{G},L)$ the enlargement of $G$ and $(G,L)$ respectively.
@@
Employing the lattice-theoretic method, we will prove the following theorem.
Theorem  1.1.
Suppose $G$ is a finite group acting symplectically and faithfully on a K3 surface $X$ in characteristic $p\geq 3$. Suppose moreover $\gcd(|G|,p)=1$. Then
@@
    If $X$ is supersingular with Artin invariant $\sigma$, then $\operatorname{rank}(S_{G}(X))\leq 22-2\sigma$. Moreover, the enlarged pair $(\widetilde{G},S_{G}(X))$ is isomorphic to a pair in Table LABEL:table:_sublist_of_HM.
@@
    If $X$ is of finite height $h$, then $\operatorname{rank}(S_{G}(X))\leq 21-2h$, and there exists a primitive embedding of $S_{G}(X)$ into the K3 lattice.
In any case, $(G,S_{G}(X))$ is a subpair of $(\operatorname{Co}_{0},\mathbb{L})$ and $\operatorname{rank}(S_{G}(X))+\ell(A_{S_{G}(X)})\leq 22$.
By Höhn–Mason’s table in [HM16], a Leech pair $(G,S)$ such that $\operatorname{rank}(S)+\ell(A_{S})\leq 22$ implies that $G$ is a subgroup of $M_{23}\leq\operatorname{Co}_{0}$. Therefore, from Theorem 1.1, any finite group of tame symplectic automorphisms of a K3 surface in odd characteristic is a subgroup of $M_{23}$. This is highly compatible with the classification by Dolgachev–Keum [DK09, Theorem 4.7] and the recent result by Ohashi–Schütt [OS24, Proposition 8.2, Theorem 8.4].
Let $X$ be supersingular with Artin invariant $\sigma$. We can extend the action of $G$ on $N_{p,\sigma}$ to $N_{p,1}$. The lattice $S_{G}(N_{p,\sigma})$ is rootless, while $S_{G}(N_{p,1})$ might contain roots. We denote by $R$ the root lattice of $S_{G}(N_{p,1})$. Apparently, $G$ acts on $R$. Let $G_{R}$ be the image of $G$ in $\operatorname{Aut}(R)$, and $G_{L}=\mathrm{Ker}(G\to\operatorname{Aut}(R))$. Then we have exact sequence:
@@
Further exploring the action of $G_{R}$ on $R$, we introduce the concept of (pseudo) p-root pairs (see §5). We summarized most of our results on supersingular K3 surfaces proved via $p$-root pairs in §5 and §6 as follows:
Theorem  1.2.
Let $X$ be a supersingular K3 surface over $k$, with Artin invariant $\sigma$. Let $G$ be a finite group acting faithfully and symplectically on $X$. Denote by $W(R)$ the Weyl group of $R$.
@@
    $G_{R}\cap W(R)$ is always a $p$-group.
@@
    If $p=11$, then the exponent of $p$ in the order of $G$ is at most $1$.
Combining both analysis on $G_{L}$ and $G_{R}$, we also obtain an upper bound (Proposition 6.8) on the exponent of $p$ in $|G|$ for a finite group $G$ of symplectic automorphisms of a K3 surface in characteristic $p=3,5,7$.
Now let us recall how our approach is related with earlier works. Kondō [Kon98] gave another proof of Mukai’s results which is lattice-theoretic and motivates our proof here. Ogus [Ogu79] (respectively Bragg–Lieblich [BL18]) proved the crystalline Torelli theorem for supersingular K3 surfaces in char $p\geq 5$ (respectively $p\geq 3$), which makes the lattice theory a powerful tool in the study of $\operatorname{Aut}^{s}(X)$. Jang [Jan16] studied the action of a symplectic automorphism on the discriminant group of the Néron–Severi group in odd characteristics which enables us to obtain Leech pairs. The key input on lattice-theoretic side is from Höhn–Mason [HM16], Jang [Jan16] and also inspired by methods developed in Laza–Zheng [LZ22].
Structure of the paper: In §2, we collect lattice-theoretic results that will be used in our paper. A key result is Proposition 2.9. In §3 we introduce supersingular K3 surfaces and recall the crystalline Torelli theorem by Ogus, Bragg and Lieblich. In particular, we give the motivation for the concept of $p$-root pairs. In §4, we prove Theorem 1.1 and discuss its relation with previous works. In §5, we developed the new machinery of (pseudo) $p$-root pairs, and in §6 we apply the new machinery to reprove an important result by Dolgachev–Keum (Theorem 1.2, (3)) in supersingular cases and give upper bounds on the wild degree of finite groups of symplectic automophisms of K3 surfaces in characteristics $p=3,5,7,11$. In Appendix A, we include the full list of Höhn–Mason pairs admitting primitive embeddings in $N_{p,1}$.
@@ 2. Lattice-Theoretical Results
In this section, we introduce some terminologies and results from lattice theory that will be used. We refer to [Nik79a], [CS99], [LZ22, Appendix A] for more details.
@@ 2.2. Finite Quadratic Forms
    A finite symmteric bilinear form consists of a finite abelian group $A$ and a symmetric $\mathbb{Z}$-bilinear map $b:A\times A\rightarrow\mathbb{Q}/\mathbb{Z}$.
@@ 2.3. Conway–Sloane Symbol
If two even integral lattices $L$ and $L^{\prime}$ have same Conway–Sloane symbol, then the discriminant forms $(A_{L},q_{L})$ and $(A_{L^{\prime}},q_{L^{\prime}})$ are isomorphic. In other words, Conway–Sloane symbol determines the discriminant form up to isomorphism.
@@ 2.4. Niemeier Lattices and Leech Pairs
For a finite group $G$ acting faithfully on a negative definite even lattice $S$, the pair $(G,S)$ is called a Leech pair, if $S$ is rootless, the induced action of $G$ on $A_{S}$ is trivial, and the action has no non-zero fixed vectors.
@@
This proposition is originally proved in [GHV12, appendix], see also [Huy16b, Proposition 2.2] and [LZ22, Proposition 3.4].
@@ 2.5. Some Lemmas on Lattice Embedding
Suppose a group $G$ acts on a lattice $L$, such that the induced action of $G$ on $A_{L}$ is trivial. Then $G$ acts trivially on $A_{S_{G}(L)}$.
@@ 2.6. Supersingular K3 Lattices
We begin with the following proposition, in which the lattice encodes the properties of the Néron-Severi lattice of a supersingular K3 surface.
@@ 3.1. Néron–Severi Lattices
It is known that $\operatorname{NS}(X)$ is an even lattice of signature $(1,\rho(X)-1)$, see [Huy16a, §17.2]. The K3 surface $X$ is called (Shioda-)supersingular if $\rho(X)=22$.
@@
[Lie13, Definition 4.14] Let $\sigma\geq 1$ be an integer and $V$ be a $2\sigma$-dimensional $\mathbb{F}_{p}$-vector space endowed with a nondegenerate and nonneutral quadratic form. A subspace $K\subset V\otimes_{\mathbb{F}_{p}}k$ is called characteristic if
@@ 3.2. Crystalline Torelli Theorem for Supersingular K3 Surfaces
Here $c_{k}$ is the composition of $N\otimes k\rightarrow\operatorname{NS}(X)\otimes k$ and $\operatorname{NS}(X)\otimes k\rightarrow H^{2}_{crys}(X/W)\otimes k\cong H^{2}_{dR}(X/k)$. The second map is induced by the crystalline Chern class map. We put $K(X,\phi)\coloneqq(id\otimes\operatorname{Fr}_{k})^{-1}(\ker(X,\phi))$.
@@
[Ogu79, Proposition 4.3][Ogu83, §1.8] Both $\ker(X,\phi)$ and $K(X,\phi)$ are characteristic subspaces of $(pN^{\vee}/pN)\otimes_{\mathbb{F}_{p}}k$.
@@
Jang [Jan16, Proposition 3.1] shows that for any $g\in\operatorname{Aut}(X)$, $g$ acts on $A_{\operatorname{NS}(X)}\otimes_{\mathbb{F}_{p}}k$ trivially if and only if $g$ is symplectic. By Proposition 3.3, $K$ is contained in

### v3

@@ Abstract.
In 2009, Dolgachev–Keum showed that finite groups of tame symplectic automorphisms of K3 surfaces in positive characteristics are subgroups of the Mathieu group of degree $23$. In this paper, we utilize lattice-theoretic methods to investigate symplectic actions of finite groups $G$ on K3 surfaces in odd characteristics. For supersingular K3 surfaces with Artin invariants at least two, we develop a new machinery called $p$-root pairs to constrain possible symplectic finite group actions (without the assumption of tameness). The concept of $p$-root pair is closely related to root systems and Weyl groups. In particular, we provide alternative proof for many results by Dolgachev–Keum and give an upper bound for the exponent of $p$ in $|G|$.
@@ 1. Introduction
K3 surfaces play an important role in many areas of mathematics and the study of their automorphism groups has a long and rich history. Nikulin [Nik79b] first systematically studied automorphisms of K3 surfaces via lattice theory. After that, in his pioneering work [Muk88], Mukai proved (using the so-called Mathieu representations) that a finite group acts faithfully and symplectically on a complex K3 surface if and only if it is a subgroup of the Mathieu group $M_{23}$ with the induced action on the set of $24$ points having at least $5$ orbits. The problem of classifying (symplectic) automorphism groups of K3 surfaces over fields of positive characteristics turns out to be more complicated. In [DK09a], Dolgachev and Keum extended Mukai’s methods to prove that if a finite group $G$ can act symplectically and faithfully on a K3 surface over an algebraically closed field of characteristic $p>0$ coprime to $|G|$, then $G$ is isomorphic to a subgroup of $M_{23}$. They also showed that if $p\geq 13$, then any symplectic and faithful action of a finite group $G$ on a K3 surface in characteristic $p$ must be tame, namely, $\gcd(|G|,p)=1$. Using the induced action on Néron–Severi lattice, Ohashi–Schütt [OS24] recently provided a complete classification of finite groups that can act on supersingular K3 surfaces with Artin invariant $1$. Combining with Dogalchev–Keum [DK09a], this provides, for any prime $p$, a classification of finite groups which admit a tame symplectic action on a K3 surface in characteristic $p$.
In this paper, we investigate finite groups acting symplectically on K3 surfaces in odd characteristic. Our approach is also based on the induced action on Néron–Severi lattice. A novelty of this work is that we uniformly deal with all K3 surfaces in odd characteristics using Leech pairs (see Definition 2.8). In particular, we develop a new machinery called $p$-root pairs to link the wild symplectic actions on supersingular K3 surfaces of arbitrary Artin invariants with Weyl group actions on root lattices. By leveraging this relation, we obtain constraints of $p$-parts of a finite group acting symplectically and faithfully on a supersingular K3 surface in odd characteristic $p$.
Before stating our main results, we fix some terminologies. Throughout this paper, let $p$ be an odd prime, and $k$ be any algebraically closed field with $\operatorname{char}(k)=p$. Let $X$ be a K3 surface in characteristic $p$ (which means that it is a K3 surface over $k$ in this paper). We denote the Néron–Severi lattice of $X$ by $\operatorname{NS}(X)$ with rank $\rho(X)$ (which is the Picard number of $X$). It is known that $\operatorname{NS}(X)$ is an even lattice with signature $(1,\rho(X)-1)$. The discriminant group of the lattice $\operatorname{NS}(X)$ is denoted by $A_{\operatorname{NS}(X)}$. We denote the automorphism group of $X$ by $\operatorname{Aut}(X)$. An automorphism $g$ of $X$ is symplectic if and only if $g$ acts identically on $H^{0}(X,\omega_{X})$. Here $\omega_{X}$ is the canonical line bundle on $X$. We denote by $\operatorname{Aut}^{s}(X)$ the group of symplectic automorphisms of $X$. Given a finite subgroup $G\leq\operatorname{Aut}^{s}(X)$, there is an induced action of $G$ on the Néron–Severi lattice $\operatorname{NS}(X)$.
@@
Suppose a group $G$ acts faithfully on a lattice $L$ such that $L^{G}=\{0\}$ and $G$ acts trivially on $A_{L}$. Then we denote by $\widetilde{G}$ the group of all automorphisms of $L$ which act trivially on $A_{L}$. We call $\widetilde{G}$ and $(\widetilde{G},L)$ the saturation of $G$ and $(G,L)$ respectively. We call $(G,L)$ saturated if $G=\widetilde{G}$.
@@
Employing the lattice-theoretic method, we prove the following proposition.
Proposition  1.1.
Fix any odd prime $p$. Let $(G,S)$ be a Leech pair with $\operatorname{rank}(S)\leq 21$. The following are equivalent:
@@
    for each supersingular K3 surface $X/k$ in characteristic $p$ with Artin invariant $\sigma$, there is an embedding $G\hookrightarrow\operatorname{Aut}^{s}(X)$, with $(G,S_{G}(X))\cong(G,S)$;
@@
    there exists a supersingular K3 surface $X/k$ in characteristic $p$ with Artin invariant $\sigma$, such that there is an embedding $G\hookrightarrow\operatorname{Aut}^{s}(X)$, with $(G,S_{G}(X))\cong(G,S)$;
  3. (3)
    there exists a primitive embedding $S\hookrightarrow N_{p,\sigma}$.
The superspecial case of Proposition 1.1 was obtained by Ohashi–Schütt [OS24]. We collect all saturated Leech pairs $(G,S)$ that admit a primitive embedding into $N_{p,1}$ (for all odd prime $p$) in Table LABEL:table:_sublist_of_HM. This table was essentially obtained by Ohashi–Schütt [OS24]. We include it in Appendix A for the reader’s convenience. The calculation is explained in the proof of Proposition 3.10. Table LABEL:table:_sublist_of_HM will be used in understanding the size of finite groups of symplectic automorphisms of supersingular K3 surfaces with arbitrary Artin invariants, see Proposition 3.12 and §6.
An action of a finite group $G$ on a K3 surface in characteristic $p$ is called tame if $\gcd(|G|,p)=1$. We prove the following two results about tame symplectic actions.
Proposition  1.2.
For a tame symplectic faithful action of a finite group $G$ on a supersingular K3 surface $X$ in characteristic $p\geq 3$ with Artin invariant $\sigma$, we have $\operatorname{rank}(S_{G}(X))\leq 22-2\sigma$. Moreover, the saturated pair $(\widetilde{G},S_{G}(X))$ is isomorphic to a pair in Table LABEL:table:_sublist_of_HM. Conversely, each pair in Table LABEL:table:_sublist_of_HM (with corresponding prime $p$) is isomorphic to $(G,S_{G}(X))$ for a tame action of a finite group $G$ on a superspecial K3 surface $X$ in characteristic $p$.
Proposition  1.3.
For a tame symplectic faithful action of a finite group $G$ on a K3 surface $X$ of finite height $h$ in characteristic $p\geq 3$, we have $\operatorname{rank}(S_{G}(X))\leq 21-2h$, and there exists a primitive embedding of $S_{G}(X)$ into the K3 lattice.
In both situation of Proposition 1.2 and Proposition 1.3, the pair $(G,S_{G}(X))$ is always a subpair of $(\operatorname{Co}_{0},\mathbb{L})$ and $\operatorname{rank}(S_{G}(X))+\ell(A_{S_{G}(X)})\leq 22$. By Höhn–Mason’s table in [HM16], a Leech pair $(G,S)$ such that $\operatorname{rank}(S)+\ell(A_{S})\leq 22$ implies that $G$ is a subgroup of $M_{23}\leq\operatorname{Co}_{0}$. Therefore, Proposition 1.2 and Proposition 1.3 provide an alternative proof (in odd characteristic cases) for the significant result by Dolgachev–Keum [DK09a, Theorem 4.7] that any finite group of tame symplectic automorphisms of a K3 surface is a subgroup of $M_{23}$. See also Ohashi–Schütt [OS24, Theorem 1.1].
Let $X$ be supersingular with Artin invariant $\sigma$. We can extend the action of $G$ on $N_{p,\sigma}$ to $N_{p,1}$. The lattice $S_{G}(N_{p,\sigma})$ is rootless, while $S_{G}(N_{p,1})$ might contain roots. We denote by $R$ the root lattice of $S_{G}(N_{p,1})$. Apparently, $G$ acts on $R$. Let $G_{R}$ be the image of $G$ in $\operatorname{Aut}(R)$, and $G_{L}=\mathrm{Ker}(G\to\operatorname{Aut}(R))$. Then we have an exact sequence:
@@
Further exploring the action of $G_{R}$ on $R$, we introduce the concept of (pseudo) p-root pairs (see §5).
Definition  1.4.
We summarize most of our results on $p$-root pairs and their applications to supersingular K3 surfaces in §5 and §6 as follows:
Theorem  1.5.
Let $(H,R)$ be a $p$-root pair, then:
@@
    $R$ cannot be of type $D_{m},m\geq 5$ (Proposition 5.11). If $R=D_{4}$, $p=3$ (Subsection 5.3).
@@
    $H\cap W(R)$ is always a $p$-group, here $W(R)$ denotes the Weyl group of $R$ (Proposition 5.23).
Proposition  1.6.
Let $X$ be a supersingular K3 surface in characteristic $p$ with Artin invariant $\sigma$. Let $G$ be a finite group acting faithfully and symplectically on $X$.
  1. (1)
    If $G$ is tame, then $G_{R}$ is trivial and $G=G_{L}$.
  2. (2)
    $(G_{R},R)$ is a $p$-root pair.
We here point out several interesting corollaries. For $p\geq 13$, we deduce that a finite group acting faithfully and symplectically on a supersingular K3 surface is always tame, see Theorem 6.5. This provides an alternative proof of [DK09a, Theorem 2.1] for supersingular K3 surfaces.
We also deduce that if $p=11$, then the exponent of $p$ in the order of $G$ is at most $1$. This corollary also follows from Dolgachev–Keum [DK09b, Theorem 1.1]. See Proposition 6.7.
Moreover, combining both analysis on $G_{L}$ and $G_{R}$, we can also obtain an upper bound on the exponent of $p$ in $|G|$ for a finite group $G$ of symplectic automorphisms of a K3 surface in characteristic $p=3,5,7$. See Proposition 6.8.
Now let us recall how our approach is related with earlier works. Kondō [Kon98] gave another proof of Mukai’s results which is lattice-theoretic and motivates our proof here. Ogus [Ogu79] (respectively Bragg–Lieblich’s unpublished work [BL18]) proved the crystalline Torelli theorem for supersingular K3 surfaces in char $p\geq 5$ (respectively $p\geq 3$), which makes the lattice theory a powerful tool in the study of $\operatorname{Aut}^{s}(X)$. Nygaard [Nyg81] and Jang [Jan16] studied the induced action of an automorphism on the Witt vector cohomology and the discriminant group of the Néron–Severi group in odd characteristics respectively, which enables us to obtain Leech pairs. The key input on lattice-theoretic side is from Höhn–Mason [HM16], Jang [Jan16] and also inspired by methods developed in Laza–Zheng [LZ22].
Structure of the paper: In §2, we collect lattice-theoretic results that will be used in our paper. A key result is Proposition 2.9. In §3 we introduce supersingular K3 surfaces and recall the crystalline Torelli theorem by Ogus, Bragg and Lieblich. In particular, we give the motivation for the concept of $p$-root pairs. In §4, we prove Proposition 1.2, 1.3 and discuss their relation with previous works. In §5, we developed the new machinery of (pseudo) $p$-root pairs, and in §6 we apply the new machinery to reprove an important result by Dolgachev–Keum (Theorem 6.5) in supersingular cases and give upper bounds on the wild degree of finite groups of symplectic automophisms of K3 surfaces in characteristics $p=3,5,7,11$. In Appendix A, we include the full list of Höhn–Mason pairs admitting primitive embeddings in $N_{p,1}$.
Acknowledgement: Bin Wang is supported by General Research Fund 14307022 and Early Career Scheme 24307121 of Michael McBreen. Zhiwei Zheng is partially supported by NSFC 12301058. We thank Prof. Looijenga for helpful discussions, especially about $p$-root pairs. We thank Matthias Schütt for many helpful comments after this paper being posted on arXiv.
@@ 2. Lattice-Theoretical Results
In this section, we introduce some terminologies and results from lattice theory that will be used. We refer to [Nik79a], [CS99, Chapter 15], [LZ22, Appendix A] for more details.
@@ 2.2. Finite Quadratic Forms
    A finite symmetric bilinear form consists of a finite abelian group $A$ and a symmetric $\mathbb{Z}$-bilinear map $b:A\times A\rightarrow\mathbb{Q}/\mathbb{Z}$.
@@ 2.3. Conway–Sloane Symbol
If two even integral lattices $L$ and $L^{\prime}$ have the same Conway–Sloane symbol, then the discriminant forms $(A_{L},q_{L})$ and $(A_{L^{\prime}},q_{L^{\prime}})$ are isomorphic. In other words, Conway–Sloane symbol determines the discriminant form up to isomorphism.
@@ 2.4. Niemeier Lattices and Leech Pairs
[LZ22, Definition 3.3] For a finite group $G$ acting faithfully on a negative definite even lattice $S$, the pair $(G,S)$ is called a Leech pair, if $S$ is rootless, the induced action of $G$ on $A_{S}$ is trivial, and the action has no non-zero fixed vectors.
@@
This proposition is originally from [GHV12, appendix], see also [MM25, Corollary 4.19] and [Zhe25].
@@ 2.5. Some Lemmas on Lattice Embedding
Suppose a group $G$ acts on an even lattice $L$, such that the induced action of $G$ on $A_{L}$ is trivial. Then $G$ acts trivially on $A_{S_{G}(L)}$.
@@ 2.6. Supersingular K3 Lattices
We begin with the following proposition, in which the lattice encodes the properties of the Néron–Severi lattice of a supersingular K3 surface.
@@ 3.1. Néron–Severi Lattices
It is known that $\operatorname{NS}(X)$ is an even lattice of signature $(1,\rho(X)-1)$, see [Huy16, §17.2]. The K3 surface $X$ is called (Shioda-)supersingular if $\rho(X)=22$.
@@
[Ogu79, Definition 3.19][Lie13, Definition 4.14] Let $\sigma\geq 1$ be an integer and $V$ be a $2\sigma$-dimensional $\mathbb{F}_{p}$-vector space endowed with a nondegenerate and nonneutral quadratic form. A subspace $K\subset V\otimes_{\mathbb{F}_{p}}k$ is called characteristic if
@@ 3.2. Crystalline Torelli Theorem for Supersingular K3 Surfaces
Here $c_{k}$ is the composition of $N\otimes k\rightarrow\operatorname{NS}(X)\otimes k$ and $\operatorname{NS}(X)\otimes k\rightarrow H^{2}_{crys}(X/W)\otimes k\cong H^{2}_{dR}(X/k)$. The second map is induced by the crystalline Chern class map. Ogus [Ogu79, Proposition 4.3] showed that $\ker(X,\phi)$ is a subspace of $(pN^{\vee}/pN)\otimes_{\mathbb{F}_{p}}k$. We then put $K(X,\phi)\coloneqq(id\otimes\operatorname{Fr}_{k})^{-1}(\ker(X,\phi))$.
@@
[Ogu83, §1.8] Both $\ker(X,\phi)$ and $K(X,\phi)$ are characteristic subspaces of $(pN^{\vee}/pN)\otimes_{\mathbb{F}_{p}}k$.
@@
Jang [Jan16, Proposition 3.1] shows that for any $g\in\operatorname{Aut}(X)$, $g$ acts on $A_{\operatorname{NS}(X)}\otimes_{\mathbb{F}_{p}}k$ trivially if and only if $g$ is symplectic. See also [Nyg81, Theorem 2.1] and [OS24, Theorem 3.1]. By Proposition 3.3, $K$ is contained in